Not in book
Write the slope-intercept form of an equation of the line that passes through (-2,0) and is parallel to the graph of y=-3x-2
Can someone help me with this one.

Found 2 solutions by checkley75, stanbon:
Answer by checkley75(3666)

y=-3x-2 the slope of this line is -3.
a line parallel to this line also has a slope of -3 & passes through (2,0) is
0=-3*2+b
0=-6+b
b=6
thus the equation of this line is
y=-3x+6

Answer by stanbon(75887)

Write the slope-intercept form of an equation of the line that passes through
(-2,0) and is parallel to the graph of y=-3x-2
---------
If the line is parallel to y=-3x-2 then m=-3.
If y=0 when x=-2, and m=-3, then
0=(-3)(2)+b
b=6
-------
EQUATION:
y=-3x+6
Cheers,
Stan H.